Circular No. 6027

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

PERIODIC COMET SHOEMAKER-LEVY 9 (1993e) P. D. Feldman, Johns Hopkins University, reports for the HST observ-

ing team the results of further analysis of FOS spectra of a region near fragment G = 15 taken on July 14: “The first two-minute segment of a 26-min exposure beginning July 14.556 UT shows strong emission of the Mg II 280-nm doublet that is not detectable in any other FOS segment (total observation time 60 min). This emission was mistakenly attributed to detector noise in the initial report of a null result (IAUC 6021). No other spectral feature is detected in the same segment. The doublet is resolved, indicating a localized source of emission in the 1′′.22 × 3′′.66 aperture, and the average brightness in the aperture is 140 Rayleighs. A threefold brightening of the continuum commences ∼ 18 min after the appearance of Mg II and is characterized by a reddening of the spectrum. The duration of this enhancement is < 20 min. At the time of the observation the fragment was 3.8×10^{6} km from Jupiter and may have crossed the Jovian magnetopause.”

Reports received via the SL9 message center describe the spectacular detection of the impact of fragment H = 14. SPIREX (cf. IAUC 6026) imaged the impact site over the limb on July 18.818 UT at 2.36 and 2.22 μm; the peak brightness appeared to be smaller than that of the G impact but still significantly larger than all other previous impacts. The Calar Alto Observing Team (cf. IAUC 6023) detected the initial brightening of the H impact on July 18.813 with a 2.3-μm filter on the 3.5-m telescope and a 3.1-μm filter on the 2.2 m telescope; the spot increased to a maximum brightness over ∼ 10 min, and impact sites D and G were still visible to the west of the H plume. K. Sekiguchi detected the H plume on July 18.814 with the South African Astronomical Observatory’s 0.75-m telescope and PtSi infrared camera at 2.2 μm; the plume became considerably brighter than that observed at the impact of fragment E and faded in ∼ 20 min. R. M. West reports that at the European Southern Observatory the H impact was first detected at 9.1–10.4 μm (3.6-m reflector + TIMMI; cf. IAUC 6026) on July 18.815, with the intensity of the plume at 10 μm ∼ 50 times that of the entire disk of Jupiter on July 18.823; it was seen without subtraction of the sky signal, indicating a temperature in excess of 300 K.

J. Lancashire, British Astronomical Association, reports his visual observation with the 0.20-m refractor at the University of Cambridge Observatories of the central part of impact site H crossing Jupiter’s central meridian on July 18.926 ± 0.003 UT, close to the prediction on IAUC 6026.

1994 July 18 Brian G. Marsden